# Neuro-Symbolic Financial Reasoning via Deterministic Fact Ledgers and Adversarial Low-Latency Hallucination Detector

### 04 Mar 2026 v1

@@ 6.3 Conclusion
In high-stakes domains such as finance, a $99\%$ accuracy rate yields $0\%$ operational trust. VeNRA establishes a novel blueprint for grounded Retrieval-Augmented Generation. By pioneering the Universal Fact Ledger (a deterministic variable store), Double-Lock Grounding (a strict extraction firewall), and the VeNRA Sentinel (a latency-optimized forensic judge), we provide a mathematically grounded architecture for reliable financial reasoning. We release the VeNRA-Data adversarial dataset and the OOM-safe Micro-Chunking Trainer to accelerate the development of specialized, logic-aware verification models across the open-source community.
Appendix A Universal Fact Ledger (UFL) 2.0 Schema
The UFL is implemented as a strict Pydantic v2 schema to ensure compatibility with the Python execution engine. It acts as the API contract between the probabilistic extractor and the deterministic solver.
@@
Appendix B Adversarial Simulation Protocols (VeNRA-Data)
@@ 3. Time Warp (Irrelevancy)
Appendix C Micro-Chunking Differential Loss Algorithm
@@
Algorithm 1 OOM-Safe Micro-Chunking Weighted Loss
@@
Appendix D VeNRA Prompting Physics (Bookending)

### 08 Mar 2026 v2

@@ 7.3 Conclusion
VeNRA establishes a novel blueprint for grounded RAG in financial reasoning by addressing the two root causes of RAG failure, retrieval conflation and arithmetic hallucination, at the architectural level rather than through post-hoc filtering. The UFL replaces probabilistic text retrieval with deterministic typed variables, and the Double-Lock Grounding protocol acts as a strict extraction firewall before they can contaminate the retrieval context. The PAL Code Agent then eliminates free-form numerical generation entirely, reducing hallucination rates to $1.2\%$ on a cross-benchmark evaluation spanning FinanceBench, FinQA, and TAT-QA.
The VeNRA Sentinel addresses the residual error surface through forensic trace auditing. Trained exclusively on programmatically constructed Ecological Errors rather than naturally occurring hallucination corpora, the 3B-parameter Sentinel achieves $91\%$ adversarial Flip Rate on the controlled evaluation— outperforming all SoTA model while operating within a ${\lesssim}150$ ms inference budget at the $28{\times}$ speedup delivered by single-token classification.
Three contributions emerge as independently transferable to other domains: the Micro-Chunking loss algorithm for training under extreme differential penalisation; the Selective Repetition prompt architecture for robust long-context claim verification; and the Adversarial Simulation data generation pipeline, which we release alongside VeNRA-Data and the Micro-Chunking Trainer to accelerate the development of logic-aware verification models across the open-source community.
Appendix A Universal Fact Ledger Schema
The UFL is implemented as a strict Pydantic schema to ensure compatibility with the Python execution engine. It acts as the API contract between the probabilistic extractor and the deterministic solver.
@@ Appendix A Universal Fact Ledger Schema
Appendix B The “Virtual Context” Protocol and HITL Refinement
To minimize context-window bloat and token costs, the Refiner operates under a “Virtual Context” protocol: it is denied access to the raw context $C$. Instead, it relies entirely on the Teacher’s $\Theta_{teacher}$ block as its “eyes.” Guided by a strict meta-prompt, the Refiner surgically corrects nomenclature confusions, sign-convention inversions (e.g., treating a cash outflow as an inflow), and boilerplate evasions, effectively overruling the Teacher model when it exhibits “math-blindness.”. If the result was still ambiguous, we tag for human annotation via the specific UI shown in Fig. 8.
Appendix C Adversarial Simulation Protocols (VeNRA-Data)
@@ 3. Time Warp (Irrelevancy)
Appendix D Micro-Chunking Differential Loss Algorithm
@@
Algorithm 1 Micro-Chunking Weighted Loss
@@
Appendix E VeNRA Prompting Physics (Bookending)